# A Black-Box Evaluation Framework for Semantic Robustness in Bird’s Eye View Detection

### v2

@@ Introduction
This work begins by formulating the semantic perturbation threat model. We notice that adopting the mean Average Precision (mAP) metric as an objective function for optimisation presents challenges due to its discontinuous nature. To address this issue, we design a distance-based surrogate objective function that incorporates the bounding box matching mechanism used in BEV detection tasks. As defined in Eq. (3), our surrogate function is continuous for matched bounding boxes while exhibiting a strong negative correlation with detection precision. These characteristics enable the function to serve as an effective proxy for optimising semantic perturbations aimed at degrading the performance of targeted BEV models.
Furthermore, we propose a query-based global optimisation algorithm called SimpleDIRECT, to solve the optimisation problem by finding the ‘worst-case’ perturbation that compromises the target BEV models. In this algorithm, we boost the optimisation performance and computational efficiency of the DIRECT optimisation (Piyavskii 1972) by reducing redundancies in identifying potential optimal solutions and leveraging slope information to guide the optimisation process. To evaluate the semantic robustness of camera-based BEV detection models, we introduce an evaluation framework focusing on three common semantic perturbations: geometric transformation, colour shift, and motion blur. These perturbations represent typical challenges posed by camera-system anomalies and real-world driving scenarios. We conduct extensive experiments on the nuScenes dataset (Caesar et al. 2020) to assess the performance of recent BEV models under these challenging conditions. We empirically demonstrate the superiority of our proposed evaluation framework over randomised perturbations used in previous literature and two optimisation-based baseline methods. In our robustness benchmark on ten BEV models, our evaluation compromised the BEVDet (Huang et al. 2021) that shows considerable resilience against natural corruption. These findings highlight shortcomings in previous literature (Zhu et al. 2023; Xie et al. 2023a) and the critical need for our work.
@@
In this work, we focus on the object detection task in BEV perception based on multi-view images. As a fundamental task in autonomous driving, the robustness of object detection has attracted notable research interest (Kong et al. 2023a, b; Xie et al. 2023a). To assess the robustness of camera-based BEV detection models, several efforts have been made to study the pixel-level and patch-level adversarial attacks (Abdelfattah et al. 2021; Park et al. 2021; Zhang et al. 2022a; Xie et al. 2023b). Unlike adversarial attacks (Carlini et al. 2019), which often exploit model vulnerabilities with unrealistic inputs, semantic perturbations (Hendrycks and Dietterich 2019) realistically mimic naturally occurring corruptions. Therefore, semantic perturbations have been widely considered in robustness evaluations to assess model performance under real-world conditions (Mirza et al. 2021; Yu et al. 2023). Most recently,  Kong et al. (2023a) and Xie et al. (2023a) introduced natural corruption respectively on LiDAR and camera-based BEV models, and Zhu et al. (2023) studied the influence of common visual corruptions. While these approaches provide valuable insights, they may not comprehensively reveal the underlying vulnerabilities of BEV models. Their limitation lies in relying on randomised perturbations to assess the semantic robustness of these models.
@@
As a specific implementation of DO, DIRECT divides the PO nodes following a space trisection strategy. Without loss of generality, consider a $d$-dimensional node $\Theta_{r}$ at depth $h$ that is poised for further exploration. $\Theta_{r}$ is shaped as a hyperrectangle characterised by $m$ dimensions each with longer sides of length $3^{-h}$ and $d-m$ dimensions each with shorter sides of length $3^{-h-1}$, where $0<m\leq d$. DIRECT performs query sampling within those $m$ dimensions that correspond to long edges. The locations of sampled points are given by $\theta_{r}\pm 3^{-h-1}\mathbf{e}_{i}$, for $i\in\{1,\ldots,m\}$, where $\mathbf{e}_{i}$ is a unit vector along $i$-th dimension. After evaluating the objective function at all newly sampled points, DIRECT equally divides each long side into three and forms new subspaces. The produced subspaces are centred at sampled points. Critically, the points yielding larger values of the objective function are allocated correspondingly larger subspaces. This principle ensures that areas of the search space with more promising results are explored more thoroughly (Jones et al. 1993). Besides, Wang  et al. (Wang et al. 2023) proposed to track the slopes between the queried points during space trisection. To achieve this, they enabled DIRECT to compute the slopes along each long edge as follows:
@@ 21      $\mathcal{P}=\text{NodeSelection()}$()
$\max_{\Theta_{q}\in\mathcal{I}^{h}_{1}}\frac{\mathcal{L}\left(\theta_{p}\right)-\mathcal{L}\left(\theta_{q}\right)}{\delta_{p}-\delta_{q}}\geq\min_{\Theta_{q}\in\mathcal{I}^{h}_{2}}\frac{\mathcal{L}\left(\theta_{q}\right)-\mathcal{L}\left(\theta_{p}\right)}{\delta_{q}-\delta_{p}};$ (5)
${\left|\mathcal{L}_{\max}\right|}\cdot\epsilon\leq\mathcal{L}\left(\theta_{p}\right)-\mathcal{L}_{\max}-\delta_{p}\min_{\Theta_{q}\in\mathcal{I}^{h}_{2}}\frac{\mathcal{L}\left(\theta_{q}\right)-\mathcal{L}\left(\theta_{p}\right)}{\delta_{q}-\delta_{p}}.$ (6)
@@
(b) The number of nodes selected by different conditions
@@ A Case Study on the Full Validation Set
Limited by the GPU resources, we could not afford to conduct the benchmark on the full validation set that contains 150 scenes (6019 frames). In this case study, we extend our method to the full validation set by perturbing selected frames and subsequently applying the perturbation to other frames. Specifically, given that each scene contains approximately 40 frames, we perturb the initial frame of a scene and apply the resulting perturbation to the following nineteen frames. We then update the perturbation at ${}^{\text{th}}$) frame and apply it to the remaining frames in the scene. We adopt the small version of BEVFormer with and without using temporal information as two target models to study the impact of temporal information on the robustness performance. Additionally, we conduct the random search with five attempts on all frames and report the best perturbation result on each frame as a baseline here to approximate the natural corruption in existing literature (Zhu et al. 2023; Xie et al. 2023a). As evidenced in Tab. 4, despite the perturbations being optimised on only a small fraction (5%) of the validation set, the proposed framework consistently outperforms randomised perturbation applied on the entire dataset. It can be observed that the model’s robustness could benefit from the temporal information. While both models could resist random perturbations, the model with temporal information has a notably smaller performance drop under our optimised perturbation, compared to when the temporal information is disabled.
@@ Conclusion
In this work, we propose a query-based black-box framework for evaluating the ‘worst-case’ robustness of BEV detection models against semantic perturbations. In our experiment, this method outperforms natural perturbation and strong optimisation-based baselines and significantly reduces the precision of camera-based BEV models, demonstrating the effectiveness of the proposed framework for measuring the vulnerability of existing BEV models. As autonomous driving technology continues to evolve, the proposed evaluation framework can be easily deployed on newly developed perception models due to its black-box nature. However, our robustness evaluation is still limited to certain semantic perturbations. Exploring how to develop more practical and representative perturbations is a promising future research direction.
@@ Acknowledgments
Figure 3: An illustration of the first three iterations of solving a 2-D problem.
In Fig. 3, we illustrate the space trisection process across the first three iterations. The algorithm identifies potential optimal subspaces for trisection during each iteration, methodically exploring the selected areas.
@@
$\min_{\Theta_{q}\in\mathcal{I}^{h}_{2}}\frac{\mathcal{L}\left(\theta_{q}\right)-\mathcal{L}\left(\theta_{p}\right)}{\delta_{q}-\delta_{p}},$

### v3

@@ Introduction
This work begins by formulating the semantic perturbation threat model. We notice that adopting the mean Average Precision (mAP) metric as an objective function for optimisation presents challenges due to its discontinuous nature. To address this issue, we design a distance-based surrogate objective function that incorporates the bounding box matching mechanism used in BEV detection tasks. As defined in Eq. (3), our surrogate function is continuous for matched bounding boxes while exhibiting a strong negative correlation with detection precision. These characteristics enable the function to serve as an effective proxy for optimising semantic perturbations aimed at degrading the performance of targeted BEV models. Furthermore, we propose a query-based global optimisation algorithm called SimpleDIRECT, to solve the optimisation problem by finding the ‘worst-case’ perturbation that compromises the target BEV models. In this algorithm, we boost the optimisation performance and computational efficiency of the DIRECT optimisation (Piyavskii 1972) by reducing redundancies in identifying potential optimal solutions and leveraging slope information to guide the optimisation process. To evaluate the semantic robustness of camera-based BEV detection models, we introduce an evaluation framework focusing on three common semantic perturbations: geometric transformation, colour shift, and motion blur. These perturbations represent typical challenges posed by camera-system anomalies and real-world driving scenarios. We conduct extensive experiments on the nuScenes dataset (Caesar et al. 2020) to assess the performance of recent BEV models under these challenging conditions and empirically demonstrate the superiority of our proposed evaluation framework over randomised perturbations used in previous literature and two optimisation-based baseline methods. In our robustness benchmark on ten BEV models, our evaluation compromised the BEVDet (Huang et al. 2021) that shows considerable resilience against natural corruption. These findings highlight shortcomings in previous literature (Zhu et al. 2023; Xie et al. 2023a) and the critical need for our work.
@@
In this work, we focus on the object detection task in BEV perception based on multi-view images. As a fundamental task in autonomous driving, the robustness of object detection has attracted notable research interest (Kong et al. 2023a, b; Xie et al. 2023a). To assess the robustness of camera-based BEV detection models, several efforts have been made to study the pixel-level and patch-level adversarial attacks (Abdelfattah et al. 2021; Park et al. 2021; Zhang et al. 2022a; Xie et al. 2023b). Unlike adversarial attacks (Carlini et al. 2019), which often exploit model vulnerabilities with unrealistic inputs, semantic perturbations (Hendrycks and Dietterich 2019) mimic naturally occurring corruptions. Therefore, semantic perturbations have been widely considered in robustness evaluations to assess model performance under real-world conditions (Mirza et al. 2021; Yu et al. 2023). Most recently,  Kong et al. (2023a) and Xie et al. (2023a) introduced natural corruption respectively on LiDAR and camera-based BEV models, and Zhu et al. (2023) studied the influence of common visual corruptions. While these approaches provide valuable insights, they may not comprehensively reveal the underlying vulnerabilities of BEV models. Their limitation lies in relying on randomised perturbations to assess the semantic robustness of these models.
@@
As a specific implementation of DO, DIRECT divides the PO nodes following a space trisection strategy. Without loss of generality, consider a $d$-dimensional node $\Theta_{r}$ at depth $h$ that is poised for further exploration. $\Theta_{r}$ is shaped as a hyperrectangle characterised by $m$ dimensions each with longer sides of length $3^{-h}$ and $d-m$ dimensions each with shorter sides of length $3^{-h-1}$, where $0<m\leq d$. DIRECT performs query sampling within those $m$ dimensions that correspond to long edges. The locations of sampled points are given by $\theta_{r}\pm 3^{-h-1}\mathbf{e}_{i}$, for $i\in\{1,\ldots,m\}$, where $\mathbf{e}_{i}$ is a unit vector along $i$-th dimension. After evaluating the objective function at all newly sampled points, DIRECT divides each long side into three equal parts, creating new subspaces. These subspaces are centred at the sampled points. Crucially, points yielding optimal results are allocated proportionally larger subspaces within $\Theta_{r}$ during partition. This principle ensures that areas of the search space with more promising results are explored more thoroughly (Jones et al. 1993). Besides, Wang et al. (2023) proposed to track the slopes between the queried points during space trisection. To achieve this, they enabled DIRECT to compute the slopes along each long edge as follows:
@@ 21      $\mathcal{P}=\text{NodeSelection()}$()
$\min_{\Theta_{q}\in\mathcal{I}^{h}_{1}}\frac{\mathcal{L}\left(\theta_{p}\right)-\mathcal{L}\left(\theta_{q}\right)}{\delta_{p}-\delta_{q}}\geq\max_{\Theta_{q}\in\mathcal{I}^{h}_{2}}\frac{\mathcal{L}\left(\theta_{q}\right)-\mathcal{L}\left(\theta_{p}\right)}{\delta_{q}-\delta_{p}};$ (5)
${\left|\mathcal{L}_{\max}\right|}\cdot\epsilon\leq\mathcal{L}\left(\theta_{p}\right)-\mathcal{L}_{\max}-\delta_{p}\max_{\Theta_{q}\in\mathcal{I}^{h}_{2}}\frac{\mathcal{L}\left(\theta_{q}\right)-\mathcal{L}\left(\theta_{p}\right)}{\delta_{q}-\delta_{p}}.$ (6)
@@
(b) Counting the PO nodes
@@ 9      $\mathcal{P}=\tilde{\vartheta}$
Figure 2: A illustration of the impact of semantic perturbations operated at different strengths on five BEV perception models. The effects are quantified regarding the number of matches (# Match) and the distance metric defined in Eq. (3). Each figure starts with clean input frames, while we set the following perturbations to $\gamma\in[0.1,0.4]$ for colour, $\gamma\in[0.04,0.1]$ for geometry, and kernel sizes ${5,7,9,11}$ for motion blur. The shaded area indicates the standard deviations of the models’ performance.
@@ A Case Study on the Full Validation Set
Limited by the GPU resources, we could not afford to conduct the benchmark on the full validation set that contains 150 scenes (6019 frames). In this case study, we extend our method to the full validation set by perturbing selected frames and subsequently applying the perturbation to other frames. Specifically, given that each scene contains approximately 40 frames, we perturb the initial frame of a scene and apply the resulting perturbation to the following nineteen frames. We then update the perturbation at ${}^{\text{th}}$) frame and apply it to the remaining frames in the scene. We adopt the small version of BEVFormer with and without using temporal information as two target models to study the impact of temporal information on the robustness performance. Additionally, we conduct the random search with five attempts on all frames and report the best perturbation result on each frame as a baseline here to approximate the natural corruption in existing literature (Zhu et al. 2023; Xie et al. 2023a). As evidenced in Tab. 4, despite the perturbations being optimised on only a small fraction (5%) of the validation set, the proposed framework consistently outperforms randomised perturbation applied on the entire dataset. While both models could resist random perturbations, the model with temporal information has a notably smaller performance drop under our optimised perturbation, compared to when the temporal information is disabled.
@@ Conclusion
In this work, we propose a query-based black-box framework for evaluating the ‘worst-case’ robustness of BEV detection models against semantic perturbations. In our experiment, this method outperforms natural perturbation and strong optimisation-based baselines and significantly reduces the precision of camera-based BEV models, demonstrating the effectiveness of the proposed framework for measuring the vulnerability of existing BEV models. As autonomous driving technology continues to evolve, the proposed evaluation framework can be easily deployed on newly developed perception models due to its black-box nature. However, our robustness evaluation is still limited to certain semantic perturbations. Developing more practical and representative perturbations offers a promising avenue for future research.
@@ Acknowledgments
Figure 3: An illustration of the first three iterations of solving a 2-D maximisation problem.
In Fig. 3, we illustrate the space trisection process across the first three iterations. The algorithm identifies potential optimal subspaces for trisection during each iteration, methodically exploring the selected areas. While the original DIRECT was proposed to solve minimisation problems, we adopted it to handle a maximisation problem, where the slopes computed in Fig. 1(c) are more likely to be negative.
@@
$\max_{\Theta_{q}\in\mathcal{I}^{h}_{2}}\frac{\mathcal{L}\left(\theta_{q}\right)-\mathcal{L}\left(\theta_{p}\right)}{\delta_{q}-\delta_{p}},$